Suppose the demand for a product is given by P = 30 – 2Q. Also, the supply is given by P = 5 + 3Q. If a $5 per-unit excise tax is levied on the buyers of a good, after the tax, the total amount of tax paid by the producers is
Answer #1

Demand: P = 30 - 2Q

Supply: P = 5 + 3Q

Before tax equilibrium price:

30 - 2Q = 5 + 3Q

Q = 5 and P = 20

New Demand:

P = (30 - 5) - 2Q

P = 25 - 2Q

Equating New Demand with Supply:

25 - 2Q = 5 + 3Q

Q = 4

P = 17 (price received by producers)

Tax paid by producers = 20 - 17 = $ 3 per unit

(Total tax paid by producers = 3 x 4 = $ 12)